12123333333333333333:

What is the perimeter, P, of a rectangle that has a length of x + 6 and a width of y − 1? P = x + y + 5 P = x + y − 7 P = 2x + 2y + 10

12123333333333333333:

@justus @mhchen

justus:

Thats correct, Good job

12123333333333333333:

Joy is training for a triathlon. She runs a distance of 2x miles, cycles a distance of (x2 + 7) miles, and swims a distance of (x − 2) miles. Which simplified expression is equivalent to the total distance, in miles, Joy covers? x2 + 3x + 5 x2 + 3x + 9 3x2 + 5 3x2 + 9

justus:

2x+x²+7+x-2=x²+2x+x+7-2= x²+3x+5=total distance

12123333333333333333:

The function f(x) = 4x + 5,000 represents the amount of money a television is being sold for, where x is the number of televisions being manufactured. The function g(x) = 20x − 500 represents the cost of production, where x is the number of televisions being manufactured. What is (f − g)(100)? Explain. $6,900 is the profit made from 100 televisions. $3,900 is the profit made from 100 televisions. $6,900 is the cost of manufacturing 100 televisions. $3,900 is the cost of manufacturing 100 televisions.

justus:

F(x)-g(x)=4x+5000-(20x-500) 4x-20x+5000+500 -16x+5500 -16*(100)+5500=3900 3900=3.9k

justus:

I believe it is B

12123333333333333333:

okey @mhchen is it B?

justus:

Im positive. Don't worry its right

12123333333333333333:

Which of the following shows that polynomials are closed under subtraction when two polynomials, (5x62 + 3x + 4) − (2x^2 + 5x − 1), are subtracted? 3x^2 − 2x + 5; will be a polynomial 3x^2 − 2x + 5; may or may not be a polynomial 3x^2 + 8x + 3; will be a polynomial 3x^2 + 8x + 3; may or may not be a polynomial sorry i am makin u help soo much.

12123333333333333333:

It is not the ones with 8s..

justus:

Its okay let me try

12123333333333333333:

It's....A? :D

justus:

Yes! Great job

12123333333333333333:

Which of the following shows that polynomials are closed under addition when two polynomials 4x2 − 8x − 7 and −5x + 16 are added? 4x2 − 13x + 9; may or may not be a polynomial 4x2 + 13x − 23; may or may not be a polynomial 4x2 − 13x + 9; will be a polynomial 4x2 + 13x − 23; will be a polynomial

justus:

What do you think?

12123333333333333333:

Hmm C? Is that correct??

justus:

Yes. Amazing :)

12123333333333333333:

Thank you

justus:

No problem!

12123333333333333333:

THANK YOU SO MUCH!!!!

12123333333333333333:

YOU ARE THE BEST!!

justus:

You're welcome!

justus:

I appreciate that

12123333333333333333:

NO. You don't get it!!

12123333333333333333:

I had a C

justus:

Oof

justus:

You put a different option by accident?

12123333333333333333:

12123333333333333333:

THANK YOUUU!! MY MOM WILL BE HAPPPYYY

justus:

You're welcome! I hope she always is

12123333333333333333:

THANK YOUU!!!!!!!!!!